Out of 54 randomly selected patients of a local hospital who were surveyed, 49 reported thatthey were satisfied with the care they received. Construct and interpret a 95% confidenceinterval for the percentage of all patients satisfied with their care at the hospital.
1
Expert's answer
2019-11-14T09:24:44-0500

Identify the sample statistics "n=54, X=49."

Find the point estimates


"\hat{p}={49 \over 54}\approx0.9074, \hat{q}\approx0.0926"

Verify that the sampling distribution of "\hat{p}" can be approximated by the normal distribution


"n\hat{p}=49>5, n\hat{q}=5\geq5"

Find the critical value "z^*" that corresponds to the given level of significance


"\alpha=0.05, z^*=z_{\alpha/2}=z_{0.05/2}=1.96"

The 95% confidence interval is given below:


"\hat{p}\pm z_{0.05/2}\sqrt{{\hat{p}(1-\hat{p}) \over n}}"

"0.9074\pm 1.96\sqrt{{0.9074(1-0.9074) \over 54}}"

"0.9074\pm 0.0773"

"(0.8301,0.9847)"

"Low: 0.8301"

"High: 0.9847"

Therefore, the 95% confidence interval for the percentage of all patients satisfied with their care at the hospital is 


"(0.8301,0.9847)"


We are "95\%" confident that the percentage of all patients satisfied with their care at the hospital is between "83.01\%"

and "98.47\%"